OpenStudy (anonymous):

the gasoline gauge on a van initially read 3/8 full. when 12 gallons were added to the tank, the gauge read 5/8 full. how many more gallons are needed to fill the tank?

OpenStudy (anonymous):

Let \(x\) be the size of the tank. Can you create an equation based on the information given?

OpenStudy (anonymous):

We could write 3/8 of tank as \(\frac 3 8 x\) and 5/8 of tank as \(\frac 5 8 x\)

OpenStudy (anonymous):

"3/8 full. when 12 gallons were added" becomes \(\frac 3 8 x + 12\)

OpenStudy (anonymous):

So our final equation is \[ \frac 3 8 x + 12 = \frac 5 8 x \]

OpenStudy (anonymous):

So 2/8 of the tank was added and that was 12 gallons. 2/8 is 1/4. A quarter of the tank is 12 then the whole tank would be 48 right? So you'll need 36 more gallons.

OpenStudy (goformit100):

@veronica192 "Welcome to OpenStudy. I can answer your questions or guide you. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."